# INQUIRE-Search: A Framework for Interactive Discovery in Large-Scale Biodiversity Databases

### v2

@@ Abstract
Large community science platforms such as iNaturalist contain hundreds of millions of biodiversity images that often capture ecological context on behaviors, interactions, phenology, and habitat. Yet most ecological workflows rely on metadata filtering or manual inspection, leaving this secondary information inaccessible at scale. We introduce INQUIRE-Search, an open-source system [^1] The codebase will be published open-source at https://github.com/Beery-Lab/INQUIRE-Search/tree/main. that enables scientists to rapidly and interactively search within an ecological image database for specific concepts using natural language, verify and export relevant observations, and utilize this discovered data for novel scientific analysis. Compared to traditional methods, INQUIRE-Search takes a fraction of the time, opening up new possibilities for scientific questions that can be explored. Through five case studies, we show the diversity of scientific applications that a tool like INQUIRE-Search can support, from seasonal variation in behavior across species to forest regrowth after wildfires. These examples demonstrate a new paradigm for interactive, efficient, and scalable scientific discovery that can begin to unlock previously inaccessible scientific value in large-scale biodiversity datasets. Finally, we emphasize using such AI-enabled discovery tools for science call for experts to reframe the priorities of the scientific process and develop novel methods for experiment design, data collection, survey effort, and uncertainty analysis.
@@ 1.1 The Need for Tools to Make Ecological Databases Discoverable
Image data have become a key component of ecological research, collected to document and analyze biodiversity, species interactions, and environmental changes across spatial and temporal scales [oliver2023camera, kays2020born]. The widespread adoption of digital photography and the deployment of automated sensors has led to a significant increase in the size and scope of ecological image databases. As an example, community science platforms like iNaturalist [iNaturalist2024] contain hundreds of millions of image-based species observations, an unprecedented scale, with tremendous potential for ecological science [callaghan2021three].
So far, the scientific use of such databases has primarily focused on documenting and analyzing species occurrences (the identification of a species at a specific place and time) [kumar2019using, humphreys2019seasonal]. But many facets of ecological study, such as biotic and abiotic interactions, resource use, and habitat associations, require information that extends beyond species identification and location alone. We refer to this contextual information as secondary data: ecologically relevant features that extend beyond the primary observation [pernat2024overcoming, marques2024retrieving, davison2025automated]. In image data, each observation shows organisms in their realized microhabitats, often with co-occurring species, observable behaviors, or signs of disturbance. Efficiently unlocking this information would enable these large image databases to realize their potential scientific utility. [beccacece2025exploring, pernat2024extracting, hu2025global].
@@
Figure 1: INQUIRE-Search pipeline overview. Once a scientific question is posed, a user can leverage our pipeline to flexibly discover the relevant data and perform analysis iteratively. Our pipeline involves the following stages: (1) Query and Prioritize to retrieve relevant images from the iNaturalist repository using natural language queries, (2) Refine and verify to manually allocate human effort and inspect top retrievals for correctness, and (3) Export and analyze to perform downstream analysis on the discovered data. This system is intended to be iterative: users can return to (1) after (2) or even (3) and repeat the process until sufficient data collection has been fulfilled.
@@ 1.2 Utilizing Machine Learning to Process Large Scale Data
Recent advances in machine learning, particularly in computer vision, provides valuable tools for dataset curation and filtering. In ecology, automated species detection and classification models are used at scale to filter large-scale databases [norouzzadeh2018automatically, willi2019identifying]. However, the supervised learning approaches to train these models require large, well-labeled datasets, which are often scarce and expensive to annotate for ecological information of interest. Moreover, scientific discovery often targets rare or context-dependent phenomena—such as predator-prey interactions, behavioral displays, or signs of disturbance—that cannot be easily captured through a fixed set of pre-defined categories. For example, modeling interactions may require defining $\binom{n}{2}$ species-pair categories for each interaction type, quickly leading to an impractically large and sparsely populated label space. This traditional closed-set design limits conventional supervised models, constraining their ability to generalize beyond predefined categories and hindering their use in open-ended ecological exploration.
Vision-language models (VLMs) offer a powerful alternative. VLMs are trained on large-scale image-text datasets, creating a joint embedding space that supports similarity-based search across modalities. Crucially, VLMs enable natural language-based querying of image databases, allowing researchers to search and analyze ecological imagery through flexible, human-readable prompts. Some of these models exhibit broad, open-world understanding of visual content and can interpret complex ecological scenes without task-specific supervision [vendrow2024inquire].
@@
The cross-modal capabilities of VLMs emerge from their training objective, which encourages the model to map both images and language into a single high-dimensional "representation space" where similar concepts are close together, and different concepts are far apart [radford2021learning]. As a simple example, an image of a hermit crab wearing plastic as a shell would be embedded into this space closer to the corresponding textual description of the scene (i.e. "a hermit crab with plastic trash as a shell"), and farther from the embedding of the text "a yellow car," as seen in Figure˜2. These embeddings capture the semantic content of images and text across many different dimensions—color, texture, shape, taxonomy, phylogeny, habitat, etc. These models therefore enable us to search for relevant images in a database through a simple similarity search—ranking each image in the database based on its distance to the embedding of a text query.
@@
Figure 2: (Top) Joint vision-language representation space. VLMs are trained to learn a joint embedding space between vision and text modalities, where similar semantic concepts are stored closer together and different concepts are farther apart. (Bottom) Image-text search using FAISS indices. At search-time, the input text query is embedded by the text encoder of the VLM. Then the query embedding vector is compared against the pre-computed iNaturalist image embeddings to retrieve the most relevant images.
@@
This type of embedding-enabled search with VLMs has been explored for ecological image databases in several recent works. WildCLIP [gabeff2024wildclip] fine-tuned a VLM with camera trap data and ecology-specific language descriptions. In training over large quantities of data, this model was able to generalize from concepts seen during training to search large wildlife datasets for new behaviors or habitat details. More recently, the INQUIRE Benchmark (INQUIRE-Bench) [vendrow2024inquire] was introduced to evaluate state-of-the-art VLMs with expert-level open-ended ecological queries across five million iNaturalist 2024 images. While the benchmark continues to pose a significant challenge for even the most capable vision-language models, it was instrumental in identifying which vision-language models perform best for these types of ecological questions. Both works focus primarily on benchmarking and improving retrieval accuracy, and clearly demonstrate the potential of such an approach to unlocking new knowledge in databases. Realizing this potential, however, requires a system that facilitates rigorous ecological inference by enabling scientists to rethink experimental design and carefully define data discovery and analysis workflows for addressing specific scientific questions.
1.3 INQUIRE-Search: a System for Interactive and Efficient Ecological Retrieval
To address this need, we introduce INQUIRE-Search, a system that enables efficient, expert-driven, iterative retrieval from large ecological image datasets. The system combines the open-world understanding of modern VLMs with question-driven methodology of ecological research: INQUIRE-Search treats large-scale dataset exploration as a process of asking, refining, and validating questions. This AI-assisted approach shifts the methodological focus from exhaustive manual labeling to strategic query design and iterative refinement, guided by both human expertise and AI retrieval results.
INQUIRE-Search is built on the principles of natural language interaction, expert-guided refinement, and efficiency. Together, these principles transform static image repositories into interactive engines for discovery. The system implements this approach through an interactive, three-stage workflow illustrated in Figure˜1:
  • Query and Prioritize: Researchers pose ecological questions in plain language (e.g. "California condor with a green "26" tag on its wing"). INQUIRE-Search retrieves and prioritizes data, ranking images by their semantic relevance to the query.
  • Refine and Verify: Scientists rapidly assess and confirm ecological validity of candidate results. This verification is flexible, allowing researchers to decide how much manual labor to allocate. Based on the results, they can also iterate by manually refining or reformulating their query to sharpen the search and ensure sufficient data collection.
  • Export and Analyze: Curated image sets can be exported with associated metadata for quantitative or qualitative analysis, supporting downstream studies or dataset curation.
INQUIRE-Search enables users to search hundreds of millions of natural-world images using natural language queries, efficiently retrieving relevant observations to minimize manual browsing. To summarize, INQUIRE-Search combines (1) a state-of-the-art vision-language embedding model that maps images and text into a shared semantic space, (2) a high-performance vector index that enables fast similarity search across hundreds of millions of embeddings, and (3) a lightweight, browser-based interface that supports filtering, inspection, verification, and data export to achieve this. Together, these components allow ecologists to iteratively query, refine, and curate datasets in real-time, transforming unstructured community science imagery into analyzable evidence for ecological research. We provide a technical overview of the pipeline and specific implementation details in the Appendix.
2 Case Studies
We worked with a diverse set of ecological experts to define a series of scientific questions that would potentially be addressable with information hidden within the iNaturalist database. We worked together to use INQUIRE-Search to address each of these questions, designing each experiment to test both the usability and utility of the tool and test the limits of what insights the system could help uncover. Each case study explores a distinct ecological question: (1) seasonal variation in bird diets, (2) post-fire forest regeneration, (3) spatio-temporal patterns of wildlife mortality, (4) plant phenology across seasonal cycles, and (5) individual Re-ID in humpback whales. These examples show how natural language-guided image search and analysis can provide a scalable and flexible approach for extracting ecological insights from large, unstructured image datasets.
2.1 Seasonal Variation in Bird Diets
2.1.1 Task Description
Diet plays a pivotal role in vertebrate life history, fundamentally influencing species’ survival, recruitment, and overall fitness. As a response trait, it is shaped by long-term environmental changes, with shifts in climate and resource availability driving dietary diversification across a wide range of taxa [burin2016omnivory, sibly2012energetics, belmaker2012global]. As an effect trait, diet provides insight into an organism’s ecological impact and its biotic interactions, reflected in its resource use and trophic position. Because of this dual role, diet has been central to advancing our understanding of key ecological and evolutionary processes, including species coexistence [abrahamczyk2015morphological], character displacement and evolutionary divergence [grant2006evolution], vertebrate diversification [burin2016omnivory, price2012tempo, rojas2012role, felice2019dietary], community assembly [lu2023food], and the impacts of global change on biodiversity [ko2014dietary].
Figure 3: Avian diets vary by species and time of year. Understanding diet enables us to connect seasonal resource use with the ecological processes that shape species interactions and community structure.
Although dietary databases exist for some vertebrate groups [wilman2014eltontraits]—including those capturing seasonal variation in diet [murphy2023SAviTraits]—they are typically compiled from published literature and often lack complete information for many species. For instance, SAviTraits 1.0 [murphy2023SAviTraits] provides seasonal dietary data for over 10,000 extant bird species, yet only about 10% of those species are recorded as exhibiting seasonal dietary variation. This might reflect not true biological invariance, but rather the absence of data on temporal shifts in diet for most species. Combined with known uncertainties in dietary classifications, it is clear that many species remain poorly studied—some may indeed exhibit seasonal dietary changes that simply have not been documented in the literature. Addressing these data gaps through traditional field-based methods would be prohibitively labor-intensive and infeasible at broad spatial scales. Current search and metadata-based filtering tools on iNaturalist are insufficient for comprehensive data extraction. For example, among more than 23,000 relevant observations, only 44 are annotated with the tag "feeding" and none are labeled "feeding on fruit." Consequently, identifying relevant cases would require manual inspection of all 23,000+ observations, which is prohibitively inefficient.
In this case study, we investigate whether INQUIRE-Search could enable systematic analysis of seasonal dietary variation in birds by searching through millions of community-sourced photographs on iNaturalist. To validate INQUIRE-Search’s performance, we compare our results against dietary information captured in SAviTraits [murphy2023SAviTraits]. We focus on five bird species with different but known diets: Gray-cheeked Thrush (Catharus minimus), Ancient Murrelet (Synthliboramphus antiquus), American Tree Sparrow (Spizelloides arborea), Red-bellied Woodpecker (Melanerpes carolinus), and American Robin (Turdus migratorius) as shown in 3. In SAviTraits, the certainty designation—derived from the number of citations supporting a species’ dietary description and the length of that description—exceeds 0.85 for these species, indicating they are among the best-studied birds in terms of diet. This makes both SAviTraits and these particular species a comprehensive and robust reference for comparison with INQUIRE-Search.
2.1.2 Methodology
We devised a systematic approach using INQUIRE-Search to extract seasonal dietary information:
@@ 2.1.2 Methodology
    For each target species, we construct natural language search queries specifying the bird species and diet type (e.g., "American robin with invertebrate in its mouth"), using the dietary classification system used in SAviTraits. This classification system includes 7 dietary categories: (1) invertebrate, (2) vertebrate, (3) seed, (4) fruit, (5) nectar and pollen, (6) carrion or animal decay, and (7) other plant matter. We applied taxonomic and temporal filters to restrict results to the target species and season (June-August for boreal summer and December-February for boreal winter), respectively. For each species, we ran 14 searches: one for each diet type-season combination.
@@
    Two researchers independently searched for and analyzed 500 images per species-diet type-season combination, each using the INQUIRE-Search tool. For each image, the researchers marked whether it contained clear dietary information for the food item corresponding to the query (i.e., the food item is visible and recognizable in the bird’s mouth).
@@
    We counted the number of marked images for each species, diet type, and season. INQUIRE-Search based dietary information was then derived from the proportion of marked images of each food type.
Figure 4: (Top) Query & Verification. Top INQUIRE-Search returns for "Red-bellied Woodpecker" and "Invertebrate" with incorrect or irrelevant images removed through expert review. (Bottom) Diet comparisons. INQUIRE-Search yields dietary patterns that align closely with those documented in the SAviTraits reference dataset.
2.1.3 Results and Analysis
Our analysis revealed substantial variation in the availability of dietary information across the five study species 4. American Robin yielded the highest number of informative (i.e., containing dietary information) images (669 total), followed by Red-bellied Woodpecker and American Tree Sparrow (395 and 188, respectively), while for both Ancient Murrelet and Gray-cheeked Thrush we found a negligible number of images containing dietary information. Among the species for which dietary information was available, seasonal patterns in image availability were also evident: American Robin had more informative images in summer than winter, while Red-bellied Woodpecker and American Tree Sparrow had more images available for winter months.
For the three species with sufficient data, dietary composition derived from INQUIRE-Search closely matched reference data from SAviTraits Figure˜4. For example, in summer, American Robin exhibited an even split between fruit and invertebrate consumption (50% each) in both datasets, while Red-bellied Woodpecker showed a diet dominated by seeds (50% in INQUIRE-Search vs. 40% in SAviTraits), with smaller contributions from invertebrates, fruits, and plant material. In winter, all three species exhibited diets strongly dominated by a single food type, and we again observed close agreement between INQUIRE-Search outputs and SAviTraits data.
2.1.4 Findings
We demonstrate that our method can effectively reproduce dietary patterns documented in established literature, provided that sufficient image data are available. The close alignment between INQUIRE-Search-derived avian diets and SAviTraits reference data highlights the potential of community science imagery to yield accurate ecological insights when properly analyzed through a systematic search methodology. However, this approach is most effective for commonly photographed species such as American Robin and Red-bellied Woodpecker. Two of the species we examined lacked sufficient image data for meaningful analysis—despite being well-studied in the scientific literature and represented in databases like SAviTraits—underscoring a fundamental constraint of community-sourced data. We anticipate this limitation will be even more pronounced for species that are genuinely understudied in the literature and for which new dietary information is most needed, as these species are also likely to be underrepresented in iNaturalist, and thus not retrievable with INQUIRE-Search. Overall, this case study shows INQUIRE-Search’s utility for rapid, large-scale analysis of visual biodiversity data, while also emphasizing its current constraints.
2.2 Post-fire Forest Regrowth
2.2.1 Task Description
Understanding how forests recover following fire is crucial for monitoring and predicting long-term ecosystem resilience. In the western US, wildfire is and has been the dominant disturbance for millennia [whitlock2008long, marlon2012long]. As a result, forests in the region are largely fire adapted and depend on wildfire for reproduction, competition reduction, and nutrient cycling [keeley2011fire]. However, the long-term consequences of fire suppression paired with recent warming and drying trends resulting from climate change are leading to wildfires that are more frequent, severe, and intense [abatzoglou2016impact, juang2022rapid]. These shifts are exceeding the adaptive capacity of many forest systems and jeopardizing their long-term resilience [harvey2016burn, davis2023reduced].
Despite the ecological importance of post-fire recovery, quantitatively assessing forest regrowth remains limited with existing survey approaches. Satellite-derived data provides broad coverage, but is generally available at spatial scales too coarse to detect individual trees and the spectral resolution of most satellites is not high enough to differentiate species [xu2021spatial]. Additionally, young trees are often spectrally indistinguishable from herbaceous species for up to 20 years after a fire [kiel2022trees]. Thus, existing satellite-based approaches struggle to capture early post-fire successional dynamics.
Field-based approaches provide valuable insights as tree seedlings are measured directly, but field campaigns are often limited in spatial coverage and temporal frequency due to cost and effort constraints. iNaturalist is a burgeoning data source that could help to overcome these constraints by providing geotagged images sourced by citizen scientists. INQUIRE-Search’s ability to efficiently search and classify iNaturalist images has the potential to reveal post-fire species and vegetation communities that are otherwise invisible to traditional remote-sensing and field-based approaches alike.
Figure 5: (Left) Forest regrowth following the 2012 High Park Fire in Colorado. With wildfires becoming more frequent and severe, fine-scale characterization of vegetation recovery is critical for understanding forest resilience. (Right) MTBS burn severity categories. We aggregate the burn categories from 30m Landsat resolution to a spatial resolution matching iNaturalist metadata (0.01 degree, $\sim$1.1km)
In this case study, we use INQUIRE-Search to identify and classify emergent forest communities growing after the 2012 High Park fire in Colorado (Figure˜5). We use INQUIRE-Search to search through nearly 20,000 ground-level iNaturalist images taken within the fire perimeter and provide a set of images containing individuals representing two target forest types: coniferous and deciduous. To test how burn severity impacts forest regeneration outcomes, we also examine the relationships between burn severity categories as classified by the Monitoring Trends in Burn Severity (MTBS) dataset (right side of Figure˜5) and each forest type.
2.2.2 Methodology
We designed a systematic approach using INQUIRE-Search to identify individuals representing emergent post-fire forest communities for each burn severity category:
  1. 1.
    To restrict images to a spatial subset containing the burn area, we bounded our search by latitude (40.57 to 40.75 degrees North) and longitude (105.18 to 105.54 degrees West) (Figure˜5). We also manually checked image timestamps to ensure all images were captured after the High Park fire.
  2. 2.
    For each target forest type, we constructed a natural language search specifying the forest type and restricting images to areas with evidence of past burning. The specific search terms are as follows:
@@ 2.2.2 Methodology
    • “young coniferous trees in burned forest”
    • “young deciduous trees in burned forest”
  3. 3.
    We exported and analyzed 200 images for each search. Using reported image coordinates, we restricted images to burned areas within the High Park MTBS fire perimeter using the sf package in R3 [pebesma2018simple].
  4. 4.

### v3

@@ Abstract
  1. 1.
    Community science platforms such as iNaturalist contain hundreds of millions of biodiversity images that capture ecological context on behaviors, interactions, phenology, and habitat. Yet ecological workflows to surface these data rely on metadata filtering or manual inspection, leaving this "secondary information" largely inaccessible at scale.
  2. 2.
    We introduce INQUIRE-Search, an open-source system that uses natural language to enable scientists to rapidly search within an ecological image database for specific concepts, verify and export relevant observations, and use these outputs for downstream scientific analysis. Compared to traditional methods, INQUIRE-Search takes a fraction of the time, opening up new possibilities for scientific questions that can be explored.
  3. 3.
    Through five case studies, we demonstrate the range of applications INQUIRE-Search can support, from seasonal variation in behavior across species to forest regrowth after wildfires. These examples demonstrate a new paradigm for interactive, efficient, and scalable scientific discovery that can begin to unlock previously inaccessible scientific value in large-scale biodiversity datasets.
  4. 4.
    Finally, we highlight how AI-enabled discovery tools for science require reframing aspects of the scientific process, including experiment design, data collection, survey effort, and uncertainty analysis.
Keywords community science, computer vision, deep learning, natural language queries, secondary data, vector search
@@ 1.1 The Need for Tools to Make Ecological Databases Discoverable
Image data have become a key component of ecological research, collected to document and analyze biodiversity, species interactions, and environmental changes across spatial and temporal scales [oliver2023camera, kays2020born]. With the widespread use of digital photography and automated sensors, community science platforms like iNaturalist [iNaturalist2024] now contain hundreds of millions of image-based species observations [callaghan2021three].
Currently, the scientific use of such databases has primarily focused on species occurrences [kumar2019using, humphreys2019seasonal], but many ecological questions, such as biotic and abiotic interactions, resource use, and habitat associations, require information that extends beyond species identification and location alone. We refer to this contextual information as secondary data: ecologically relevant features that extend beyond the primary observation [pernat2024overcoming, marques2024retrieving, davison2025automated]. Images frequently capture organisms in their microhabitats, with co-occurring species, observable behaviors, or signs of disturbance.
@@
Figure 1: INQUIRE-Search pipeline overview. Users begin with a scientific question and iteratively (1) query and prioritize relevant iNaturalist images, (2) refine and verify results with expert inspection, and (3) export the curated data for downstream analysis. The process supports returning to earlier stages as needed until sufficient data are collected.
@@ 1.2 Utilizing Machine Learning to Process Large Scale Data
Recent advances in computer vision provide valuable tools for dataset curation and filtering, and automated species detection models are widely used in ecology [norouzzadeh2018automatically, willi2019identifying]. However, supervised learning approaches to train these models require large, well-labeled datasets, which are often hard to acquire for ecological information of interest. This closed-set design limits the ability of supervised models to generalize beyond fixed categories and constrains their use in open-ended ecological exploration.
@@
Vision-language models (VLMs) offer a powerful alternative. Trained on large image-text datasets, these models map images and language into a shared high-dimensional representation space where semantically similar concepts lie close together [radford2021learning]. For example, an image of a hermit crab wearing plastic as a shell would be embedded into this space closer to the corresponding text "a hermit crab with plastic trash as a shell," and far from unrelated descriptions such as "a yellow car," as seen in Fig 2. VLMs therefore enable us to search for relevant images in a database through a simple similarity search—ranking each image in the database based on its distance to the embedding of a text query.
@@
Figure 2: (Top) Joint vision-language representation space. VLMs are trained to learn a joint embedding space between vision and text modalities, where similar semantic concepts are stored closer together. (Bottom) Image-text search. At search-time, the input text query is embedded by the text encoder of the VLM. Then the query embedding vector is compared against the pre-computed iNaturalist image embeddings to retrieve the most relevant images.
@@
Embedding-enabled search with VLMs has already shown promise for ecological applications. WildCLIP [gabeff2024wildclip] fine-tuned a VLM on camera trap images and ecology-specific language descriptions. From training on large data, this model was able to generalize from concepts seen during training to search large wildlife datasets for new behaviors or habitat details. The INQUIRE Benchmark [vendrow2024inquire] evaluates state-of-the-art VLMs on expert-level open-ended ecological queries across five million iNaturalist 2024 images. Both works clearly demonstrate the potential of such an approach to unlocking new knowledge in databases. Realizing this potential requires a system that connects the broad capabilities of VLMs with the practical needs of ecological research, enabling scientists to make effective use of large image repositories.
To address the need, we introduce INQUIRE-Search, a system for efficient, expert-driven retrieval from large ecological image datasets. As illustrated in Fig 1, the system facilitates an iterative cycle in which researchers pose questions in plain language, retrieve semantically relevant images, assess and refine those results based on ecological validity, and ultimately export curated observations for downstream analysis.
2 Materials and Methods
2.1 INQUIRE-Search System Architecture
INQUIRE-Search is an open-source system for interactive, expert-driven retrieval from large ecological image datasets. It integrates a vision–language embedding model, a high-performance vector index, and a lightweight, browser-based interface to support efficient natural-language search, verification, and data export.
The system uses a state-of-the-art Vision–Language Model (VLM), SigLIP-So400m-384-14 [zhai2023sigmoid], to embed both images and natural-language queries into a shared high-dimensional semantic space. All iNaturalist [iNaturalist2024] images were preprocessed to a uniform format and embedded using the SigLIP image encoder. These embeddings were stored in a FAISS (Facebook AI Similarity Search) [douze2025faiss] index, enabling sub-second similarity search across approximately 300 million images.
At search time, a user-provided text query is embedded and used to rank images in the FAISS index by cosine similarity. The browser-based interface supports metadata filtering (taxonomy, geography, and date), rapid inspection and expert verification of retrieved images, and export of curated observations with complete iNaturalist metadata. Verified image sets can be exported as CSV files, including observation IDs, coordinates, timestamps, taxonomic information, and file URLs, for downstream analysis. Full implementation details are provided in the Appendix.
2.2 INQUIRE-Search Pipeline Overview
We follow a standardized human-in-the-loop workflow using INQUIRE-Search.
@@ 2.2 INQUIRE-Search Pipeline Overview
    Query and Prioritize: We composed natural-language queries tailored to each ecological question and applied optional metadata filters (species, date, location). All prompts are listed in Table 1, with complete information on the filters, number of retrievals inspected, and number of samples marked for analysis provided in the Appendix.
@@
    Retrieve and Verify: For each query, the system returned top-ranked images. Depending on the case study, we screened between 200 and 500 top retrievals for relevant data. A reviewer examined each image and labeled it informative only if it met the case-specific criteria.
@@
    Export and Analyze: Informative images and their metadata were exported as CSV files for downstream analysis. Analytical procedures were specific to each case study and are described in Section 2.3.
Case study INQUIRE-Search Prompt
Seasonal variation in bird diets Prompts of the form "<species> with <diet type> in its mouth" using: Species (common names): Gray-cheeked Thrush, Ancient Murrelet, American Tree Sparrow, Red-bellied Woodpecker, American Robin Diet types: invertebrate, vertebrate, seed, fruit, nectar and pollen, carrion or animal decay, other plant matter Total: 5 species × 7 diet types = 35 prompts.
@@
Post-fire forest regrowth "Young coniferous trees in burned forest"
"Young deciduous trees in burned forest".
Wildlife mortality "Dead bird"
Plant Phenology "Milkweed germinating or emerging"
"Milkweed flowering"
"Milkweed producing seeds"
"Milkweed dying or withering / in senescence".
Whale re-identification "White underside of humpback whale fluke".
Table 1: Prompts used for retrieval across the five INQUIRE-Search case studies.
2.3 INQUIRE-Search Case Studies
We collaborated with a diverse set of ecological experts to define a series of scientific questions that would potentially be addressable with information hidden within the iNaturalist database. Together, we used INQUIRE-Search to address each of these questions, designing each experiment to test both the usability and utility of the tool and test the limits of what insights the system could help uncover. Each case study explores a distinct ecological question: (1) seasonal variation in bird diets, (2) post-fire forest regeneration, (3) spatio-temporal patterns of wildlife mortality, (4) plant phenology across seasonal cycles, and (5) individual Re-ID in humpback whales. These examples show how natural language-guided image search and analysis can provide a scalable and flexible approach for extracting ecological insights from large, unstructured image datasets.
2.3.1 CS1: Seasonal Variation in Bird Diets
Overview Diet is a fundamental response and effect trait that shapes vertebrate survival, fitness, trophic position, and ecological interactions, and it varies widely across species and environmental conditions [burin2016omnivory, sibly2012energetics, belmaker2012global]. Although large dietary databases exist [wilman2014eltontraits], seasonal diet information remains sparse: in SAviTraits 1.0 [murphy2023SAviTraits], only about 10% of more than 10,000 bird species are recorded as exhibiting seasonal dietary shifts, a pattern likely driven by data gaps rather than true biological invariance. Extracting such information from community-science imagery is also nontrivial: among over 23,000 relevant iNaturalist observations of American Robin, only 44 are tagged “feeding,” and none specify the type of food consumed. This case study evaluates whether INQUIRE-Search can systematically recover feeding events across seasons from community-sourced photographs and reproduce known dietary patterns. We focus on five species – Gray-cheeked Thrush (Catharus minimus), Ancient Murrelet (Synthliboramphus antiquus), American Tree Sparrow (Spizelloides arborea), Red-bellied Woodpecker (Melanerpes carolinus), and American Robin (Turdus migratorius – which have high dietary certainty scores in SAviTraits [murphy2023SAviTraits], providing a robust reference for comparison.
Search Strategy: For each target species, we construct natural language search queries specifying the bird species and diet type (e.g., "American robin with invertebrate in its mouth"), using the dietary classification system used in SAviTraits. This classification system includes seven dietary categories: (1) invertebrate, (2) vertebrate, (3) seed, (4) fruit, (5) nectar and pollen, (6) carrion or animal decay, and (7) other plant matter. Each diet category was queried separately for summer (June–August) and winter (December–February). Taxonomic filters restricted results to the target species.
Verification: An image was considered informative only when a clearly identifiable food item was visible in the bird’s bill. Top 500 images were screened by a reviewer for every species-diet type-season combination.
Analysis: For each combination, we counted the marked images and calculated the proportion of informative images belonging to each diet category. These proportions were then compared qualitatively with published seasonal diet compositions from SAviTraits.
2.3.2 CS2: Post-fire Forest Regrowth
Figure 3: (Left) Forest regrowth following the 2012 High Park Fire in Colorado. With wildfires becoming more frequent and severe, fine-scale characterization of vegetation recovery is critical for understanding forest resilience. (Right) MTBS burn severity categories. We aggregate the burn categories from 30m Landsat resolution to a spatial resolution matching iNaturalist metadata (0.01, $\sim$1100m).
Overview: Understanding how forests recover following fire is essential for assessing long-term ecosystem resilience, particularly in the western United States where wildfire has been the dominant disturbance for millennia [whitlock2008long, marlon2012long] and many forests are fire-adapted [keeley2011fire]. Recent warming and drying trends associated with climate change, coupled with the legacy of fire suppression, are driving fires that are more frequent, severe, and intense [abatzoglou2016impact, juang2022rapid], exceeding the adaptive capacity of many forest systems by limiting post-fire regeneration and altering species composition [harvey2016burn, davis2023reduced]. Despite the ecological importance of post-fire recovery for understanding forest resilience, successional trajectories, and long-term ecosystem change, existing approaches provide limited capacity to assess fine-scale regrowth: satellite data often lack the spatial or spectral resolution needed to detect young trees or differentiate species [xu2021spatial, kiel2022trees], while field surveys are constrained by cost and limited spatial coverage. Because community-sourced photographs can reveal early regeneration that is invisible to both satellites and sparse field campaigns, this case study evaluates whether INQUIRE-Search can identify young coniferous and deciduous individuals within the 2012 High Park Fire perimeter and examine how their occurrence varies across burn severity classes (Fig. 3).
Search Strategy: We targeted two functional groups (young coniferous trees and young deciduous trees), using text queries such as "young coniferous trees in burned forest" and "young deciduous trees in burned forest". Spatial filters restricted observations to the High Park Fire perimeter: latitude (40.57 to 40.75 N) and longitude (105.18 to 105.54 W) (Fig 3). Temporal filters restricted images to post-fire dates. After retrieval, we applied a finer spatial filter using reported image coordinates, retaining only images within the High Park MTBS fire perimeter via the sf package in R3 [pebesma2018simple].